A baseball is hit such that it travel straight upward after being struck by the bat.a fam observes that it requires 3.00s for the ball to reach it's maximum height.find(a)it initial velocity and (b) it's maximum height ignore the effect of air residence
1
Expert's answer
2020-10-19T13:22:00-0400

a)


"v=gt=(9.8)(3)=29\frac{m}{s}"

b)


"h=0.5gt^2=0.5(9.8)(3)^2=44\ m"